California allows patients to have 8 ounces of dried processed marijuana per day, which amounts to around 304 mg of Tetrahydrocannabinol (THC). On average, the amount of THC in your body will decrease at a rate of 48% every 10 days.
Write an equation that describes the amount of THC in a person's body x days after consuming 8 ounces of marijuana, if they do not consume any more THC after that.
Struggling on this problem...
Thank you in advance!

Answer by Boreal(15235)

A=304(1-.48)^x where x=10 days
A=304(0.52)^(x/10) where x= number of days
so after 1 day there is 284.76 mg
after 30 days there is 42.74 mg